# Enhancing LLMs for Power System Simulations: A Feedback-driven Multi-agent Framework

### v1

@@ Abstract
The integration of experimental technologies with large language models (LLMs) is transforming scientific research, positioning AI as a versatile research assistant rather than a mere problem-solving tool. In the field of power systems, however, managing simulations — one of the essential experimental technologies — remains a challenge for LLMs due to their limited domain-specific knowledge, restricted reasoning capabilities, and imprecise handling of simulation parameters. To address these limitations, we propose a feedback-driven, multi-agent framework that incorporates three proposed modules: an enhanced retrieval-augmented generation (RAG) module, an improved reasoning module, and a dynamic environmental acting module with an error-feedback mechanism. Validated on 69 diverse tasks from Daline and MATPOWER, this framework achieves success rates of 93.13% and 96.85%, respectively, significantly outperforming the latest LLMs (ChatGPT 4o and o1-preview), which achieved a 27.77% success rate on standard simulation tasks and 0% on complex tasks. Additionally, our framework also supports rapid, cost-effective task execution, completing each simulation in approximately 30 seconds at an average cost of 0.014 USD for tokens. Overall, this adaptable framework lays a foundation for developing intelligent LLM-based assistants for human researchers, facilitating power system research and beyond.
@@ I Introduction
Combining laboratory automation technologies with large language models (LLMs) enables automated execution of scientific experiments [1]. Related advances span the fields of mathematics, chemistry, and clinical research, including mathematical algorithm evolution [2], geometry theorem proving [3], chemical experiment design and execution [1], as well as the development and validation of machine learning approaches for clinical studies [4]. These recent achievements signal a new research paradigm, positioning AI as a research assistant for humans with natural language communication abilities, rather than merely a specialized problem solver as in the past. Establishing LLMs as research assistants also holds significant potential for advancing power systems research.
Given the heavy reliance on simulations in power systems research, developing LLM-based assistants in this field requires equipping LLMs with the capability to conduct power system simulations. Enabling LLMs to execute simulation tasks has multiple implications: (i) At the assistant level, LLMs capable of conducting simulations would allow researchers to focus more on idea-intensive activities, such as simulation design, rather than on labor-intensive tasks like simulation implementation. (ii) At the interface level, LLMs conducting simulations might offer a natural-language interface. This interface can connect simulation tasks with other upstream/downstream power system tasks using natural language as the input/output. This is particularly helpful when these tasks’ original inputs and outputs are heterogeneous (e.g., different modalities) and originally challenging to program cohesively using regular codes. (iii) At the coding level, LLMs executing simulations might be a step toward natural language coding in power systems. This might signify an evolution in programming, bringing it closer to a more intuitive, language-driven approach, a long-standing goal of programming development for decades.
However, LLMs inherently lack the capability to perform power system simulations. For recently developed simulation tools not included in LLM pre-training datasets, LLMs generally cannot execute these simulations accurately. Even for well-established tools included in pre-training data, simulation precision remains unsatisfactory. For instance, GPT-4 often has difficulty creating small distribution grids using OpenDSS [5] or writing code for simple (optimal) power flow problems [6], even though information about both OpenDSS and (optimal) power flow is available within GPT-4’s pre-training dataset. While the underlying causes of this issue have not been widely discussed and recognized in the energy domain, we propose the following factors as potential explanations:
@@
  • Complexity: The multi-step reasoning required by complex power system simulations is inherently challenging, particularly given sparse or ambiguous representations in the model’s learned patterns w.r.t. power system simulations.
@@
Figure 1: The feedback-driven multi-agent framework. It consists of an enhanced RAG module, an advanced reasoning module, and an environmental interaction module, all interconnected through an error feedback mechanism. This framework enables iterative refinement by incorporating simulation-specific knowledge, improving reasoning for complex simulation tasks, and facilitating environmental interaction to generate accurate simulation results.
@@
Despite the various applications mentioned above, only a few studies have focused directly on using LLMs for power system simulations. These studies, however, only focus on conceptualizing the potential of LLMs in the simulation field [17], showcasing their current capabilities [18, 17], and assessing their effectiveness in generating general-purpose code for power system studies [5, 6]. While these studies offer valuable insights, they did not address the limitations pointed out earlier that limit LLMs’ simulation performance. Although the standard RAG approach used in [19, 17, 6] can indeed enable LLMs to incorporate external power systems knowledge, it is unsuitable for simulation tasks. This is because the standard RAG retrieves information based on the entire request as a single unit, which fails to capture the nuanced structure of complex simulation requests, often conflating distinct function-related and option-related elements, thereby leading to inefficiencies and reduced retrieval accuracy. Consequently, existing works fall short of systematically developing and advancing LLMs’ capability to handle complex power system simulations.
@@
To bridge this gap and enhance LLMs’ capability in power system simulations, this paper proposes a modular, feedback-driven, multi-agent framework that integrates several innovative strategies, as shown in Fig. 1. Accordingly, this paper contributes in the following ways:
@@
  • We propose an enhanced RAG module with an adaptive query planning strategy and a triple-based structure (i.e. linking options, functions, and their dependencies) for the knowledge base. This module not only expands the LLM’s accessible knowledge in an efficient and cost-effective manner, but also enables LLMs to better identify and interpret simulation functions, options, and their logical relationships than the standard RAG.
@@
  • We develop an enhanced reasoning module by leveraging simulation-specific expertise, chain-of-thought prompting (CoT) and few-shot prompting. This module enables LLMs to fully understand their role, assigned tasks, reasoning pathways, and contextual knowledge (including retrieved information) in simulation tasks, thereby strengthening their reasoning capabilities when generating simulation code.
@@
  • We further propose a feedback-driven, multi-agent framework that integrates the enhanced RAG and reasoning modules with an environmental interaction and error-correction mechanism. This framework facilitates both action execution and feedback reception, providing responsive error signals to initiate adaptive adjustments for the RAG and reasoning modules to automatically correct errors, thereby enhancing the reliability of simulation outcomes.
@@
  • Through testing across various strategies, simulation environments, and a diverse range of tasks, we reveal that even the latest LLM, o1-preview, struggles with power system simulation tasks, including those involving well-established tools like MATPOWER, despite prior exposure in the pre-training of the LLM. We further reveal that high simulation success rates depend on the cumulative effect of multiple strategies. Following this idea, our framework demonstrates high success rates, enabling cost-effective, rapid task completion, thereby providing a scalable tool for power system researchers.
@@
This paper, as a substantial extension of our preliminary work in [20], is structured as follows: Section II introduces the enhanced RAG module. Section III presents the enhanced reasoning module, and Section IV describes the environmental acting module with the feedback mechanism. Finally, Section V presents case study results, while Section VI concludes the paper with key findings and future outlook.
@@ II Enhanced RAG Module
To this end, we propose an enhanced RAG module. This is specifically designed to integrate power system simulation knowledge into LLMs and reduce hallucinations. This module emphasizes the identification of essential keywords in simulation requests to facilitate more precise knowledge retrieval than the standard RAG. It includes two main components: (i) an adaptive query planning strategy, and (ii) a triple-based structure design for the knowledge base. Together, these components provide an enhanced RAG for complex power system simulation tasks.
Figure 3: General RAG diagram, including the process of external knowledge chunking, text embedding, and parallel retrieval within a vector database to produce relevant retrieval output based on the input queries and the external knowledge base. The text embedding model used in this study is from [here].
@@ II-A Adaptive Query Planning
This section addresses the question of what types of queries should be used for retrieval. In the standard RAG approach, the entire simulation request is processed as a single unit, which, as discussed (and will be demonstrated in case studies), conflates distinct elements in the request, leading to inefficiencies and reduced retrieval accuracy. In fact, simulation requests typically contain two critical elements: the functions to be used and the options to be set. Thus, we propose using functions and options as distinct retrieval queries. However, these elements are rarely stated explicitly in simulation requests; instead, they are embedded in natural language descriptions.
To address this, we develop an agent-driven adaptive query planning strategy, as shown in Fig. 2(a), which automatically extracts function-related and option-related queries from the broader request to serve as retrieval keywords. The strategy operates in two phases: semantic recognition and keyword mapping, carried out by a retrieval agent (e.g., a general LLM). In the semantic recognition phase, the agent categorizes the request into two query types: function-related and option-related. Each function-related query is then decomposed into sub-queries, each corresponding to a potential simulation function to be used. Similarly, option-related queries are broken down into sub-queries, each linked to a potential option to be configured. This systematic separation ensures independent processing of each component within the request.
@@
To enable the retrieval agent to perform both semantic recognition and keyword mapping effectively, we design a structured, general action prompt that integrates chain-of-thought prompting (CoT) [21] and few-shot prompting [22] (i.e., few-shot CoT), as depicted in Fig. 2(b). Only the few-shot examples are tool-dependent, making them modular and easily adaptable, while the rest of the prompt remains general and independent of specific simulation tools.
@@ II-B Triple-based Structure Design for Knowledge Base
In this section, we address the question of which knowledge base should serve as the retrieval repository. While each power system simulation tool includes a user manual with detailed instructions, this manual is not an ideal retrieval repository. The reasons are twofold: (i) User manuals are designed for human readability rather than automated retrieval; although readable, they are unstructured and inefficient for machine-driven queries, especially when manuals primarily consist of formulas, tables, and figures. (ii) The main challenge for LLMs in generating simulation code is understanding the logical dependencies between options and functions, as many options are function-dependent. Using only the user manual for retrieval fails to capture these complex relationships effectively.
To overcome these issues, we propose an additional, easy-to-construct retrieval repository: a triple-based structured option document. In this document, each line represents an option, providing the following structured information in sequence: (i) option name, (ii) default value/format, (iii) function dependencies, and (iv) option description. The inclusion of triples in (iii) — linking each option, its related functions, and their dependency — enables retrieval for logical relationships. As will demonstrated in case studies, this supplementary repository significantly enhances retrieval efficiency and improves the accuracy of simulation code generation by preserving the logical context.
@@ III Enhanced Reasoning Module
Even though the enhanced RAG module provides LLMs with retrieval results tailored to a simulation request, it still remains essential to strengthen the LLM’s reasoning abilities to generate correct simulation codes based on the retrieval results. This requires a coding agent (i.e., another LLM) to write codes for simulations tasks. This agent needs to fully understand its role, assigned tasks, reasoning path, and contextual knowledge, including retrieval results, when handling simulation tasks.
To address this, we propose an enhanced reasoning module, as detailed in Fig. 4. It provides structured guidance, sequential reasoning steps, and contextual knowledge to support accurate code generation by the coding agent. Details are as follows.
@@ III-B Reasoning Framework
To enable systematic, tool-independent reasoning, we develop a few-shot CoT framework, which breaks down the simulation task into the following universal actions:
@@ III-C Knowledge Integration
  • Dynamic Retrieval Knowledge: The prompt includes placeholders for dynamic retrieval results from the enhanced RAG module, allowing the agent to incorporate request-specific knowledge into its reasoning. This knowledge provides detailed information on function dependencies, option formats, and other contextual elements needed for accurate code generation.
Eventually, by integrating both static and dynamic knowledge, as well as the above structured reasoning framework, the agent is expected to generate accurate code to address the simulation request.

### v2

@@ Abstract
The integration of experimental technologies with large language models (LLMs) is transforming scientific research. It positions AI as a versatile research assistant rather than a mere problem-solving tool. In the field of power systems, however, managing simulations — one of the essential experimental technologies — remains a challenge for LLMs due to their limited domain-specific knowledge, restricted reasoning capabilities, and imprecise handling of simulation parameters. To address these limitations, this paper proposes a feedback-driven, multi-agent framework. It incorporates three proposed modules: an enhanced retrieval-augmented generation (RAG) module, an improved reasoning module, and a dynamic environmental acting module with an error-feedback mechanism. Validated on 69 diverse tasks from Daline and MATPOWER, this framework achieves success rates of 93.13% and 96.85%, respectively. It significantly outperforms ChatGPT 4o, o1-preview, and the fine-tuned GPT4o, which all achieved a success rate lower than 30% on complex tasks. Additionally, the proposed framework also supports rapid, cost-effective task execution, completing each simulation in approximately 30 seconds at an average cost of 0.014 USD for tokens. Overall, this adaptable framework lays a foundation for developing intelligent LLM-based assistants for human researchers, facilitating power system research and beyond.
@@ I Introduction
Combining laboratory automation technologies with large language models (LLMs) enables automated execution of scientific experiments [1]. Related advances span the fields of mathematics, chemistry, and clinical research, including mathematical algorithm evolution [2], geometry theorem proving [3], chemical experiment design and execution [1], as well as the development and validation of machine learning approaches for clinical studies [4]. These recent achievements signal a new research paradigm. That is, positioning AI as a research assistant for humans with natural language communication abilities, rather than merely a specialized problem solver as in the past. Establishing LLMs as research assistants also holds significant potential for advancing power systems research.
Power systems research heavily relies on simulations. To develop LLM-based research assistants in this field, LLMs must be equipped with the capability to conduct power system simulations. Enabling LLMs to execute simulation tasks has multiple implications: (i) At the assistant level, LLMs capable of conducting simulations would allow researchers to focus more on idea-intensive activities, such as simulation design, rather than on labor-intensive tasks like simulation implementation. (ii) At the interface level, LLMs conducting simulations might offer a natural-language interface. This interface can connect simulation tasks with other upstream/downstream power system tasks using natural language as the input/output. This is particularly helpful when the original inputs and outputs of these tasks are heterogeneous (e.g., different modalities), which are originally challenging to program cohesively using regular codes. (iii) At the coding level, LLMs executing simulations might be a step toward natural language coding in power systems. This might signify an evolution in programming, bringing it closer to a more intuitive, language-driven approach, a long-standing goal of programming development for decades.
However, LLMs inherently lack the capability to perform power system simulations. For recently developed simulation tools not included in LLM pre-training datasets, LLMs generally cannot execute these simulations accurately. Even for well-established tools included in pre-training data, simulation precision remains unsatisfactory. For instance, GPT-4 often has difficulty creating small distribution grids using OpenDSS [5] or writing code for simple (optimal) power flow problems [6], even though information about both OpenDSS and (optimal) power flow is available within GPT-4’s pre-training dataset. While the underlying causes of this issue have not been widely discussed and recognized in the energy domain, the following factors have been proposed as potential explanations:
@@
  • Complexity: Complex power system simulations require multi-step reasoning, which is inherently challenging. This difficulty increases when the model’s learned patterns contain sparse or ambiguous representations related to power system simulations.
@@
Figure 1: The feedback-driven multi-agent framework. It consists of an enhanced RAG module, an advanced reasoning module, and an environmental acting module, all interconnected through an error-feedback mechanism. This framework enables iterative refinement by incorporating simulation-specific knowledge, improving reasoning for complex simulation tasks, and facilitating environmental interaction to generate accurate simulation results.
@@
Despite the various applications mentioned above, only a few studies have focused directly on using LLMs for power system simulations. These studies, however, only focus on conceptualizing the potential of LLMs in the simulation field [17], showcasing their current capabilities [18, 17], and assessing their effectiveness in generating general-purpose code for power system studies [5, 6]. While these studies offer valuable insights, they did not address the limitations pointed out earlier that limit LLMs’ simulation performance. Although the standard RAG approach used in [19, 17, 6] can indeed enable LLMs to incorporate external power systems knowledge, it is unsuitable for simulation tasks. This happens because the standard RAG retrieves information based on the entire request as a single unit. As a result, it fails to capture the nuanced structure of complex simulation requests. It often conflates distinct function-related and option-related elements, leading to inefficiencies and reduced retrieval accuracy. Consequently, existing works fall short of systematically developing and advancing LLMs’ capability to handle complex power system simulations.
@@
To bridge this gap and enhance LLMs’ capability in power system simulations, this paper proposes a modular, feedback-driven, multi-agent framework. It integrates several innovative strategies, as shown in Fig. 1. Accordingly, this paper contributes in the following ways:
@@
  • It proposes an enhanced RAG module with an adaptive query planning strategy and a triple-based structure (i.e. linking options, functions, and their dependencies) for the knowledge base. This module expands the LLM’s accessible knowledge in an efficient and cost-effective manner. Also, it enables LLMs to better identify and interpret simulation functions, options, and their logical relationships than the standard RAG.
@@
  • It develops an enhanced reasoning module by leveraging simulation-specific expertise, chain-of-thought prompting (CoT) and few-shot prompting. This module enables LLMs to fully understand their role, assigned tasks, reasoning pathways, and contextual knowledge (including retrieved information) in simulation tasks. Therefore, it strengthens their reasoning capabilities when generating simulation code.
@@
  • It further proposes a feedback-driven, multi-agent framework. It integrates the enhanced RAG and reasoning modules with an environmental interaction and error-correction mechanism. This framework facilitates both action execution and feedback reception. It provides responsive error signals to initiate adaptive adjustments for the RAG and reasoning modules to automatically correct errors, thereby enhancing the reliability of simulation outcomes.
@@
  • Through testing across various strategies, simulation environments, and a diverse range of tasks, this paper reveals that even the latest LLM, o1-preview, struggles with power system simulation tasks, including those involving well-established tools like MATPOWER, despite prior exposure in the pre-training of the LLM. This paper further reveals that high simulation success rates depend on the cumulative effect of multiple strategies. Following this idea, the proposed framework demonstrates high success rates, enabling cost-effective, rapid task completion. It, therefore, provides a scalable tool for power system researchers.
Figure 3: General RAG diagram, including the process of external knowledge chunking, text embedding, and parallel retrieval within a vector database to produce relevant retrieval output based on the input queries and the external knowledge base. Algorithm 1 details the configuration settings and process for the embedding used in this paper.
Input:
datasetPath: Paths to external documents
chunkSize: Words per chunk
embeddingModel: Chosen text embedding model
apiKey: Credential for embedding
Output:
vectorDB: Vector database for documents
1
// Step 1: Chunk;
2 chunks $\leftarrow$ splitIntoChunks(datasetPath, chunkSize)
// Step 2: Embedding;
3 vectors $\leftarrow$ embedChunks(chunks, embeddingModel, apiKey)
// Step 3: Store in Vector DB;
4 vectorDB $\leftarrow$ storeVectors(vectors, apiKey)
Algorithm 1 Embedding for External Documents
@@
This paper, as a substantial extension of the authors’ preliminary work in [20], is structured as follows: Section II introduces the enhanced RAG module. Section III presents the enhanced reasoning module, and Section IV describes the environmental acting module with the feedback mechanism. Finally, Section V presents case study results, while Section VI concludes the paper with key findings and future outlook.
@@ II Enhanced RAG Module
To this end, this paper proposes an enhanced RAG module. This is specifically designed to integrate power system simulation knowledge into LLMs and reduce hallucinations. This module emphasizes the identification of essential keywords in simulation requests to facilitate more precise knowledge retrieval than the standard RAG. It includes two main components: (i) an adaptive query planning strategy, and (ii) a triple-based structure design for the knowledge base. Together, these components provide an enhanced RAG for complex power system simulation tasks.
@@ II-A Adaptive Query Planning
This section addresses the question of what types of queries should be used for retrieval. In the standard RAG approach, the entire simulation request is processed as a single unit. As discussed and demonstrated in case studies, this conflates distinct elements in the request. As a result, it leads to inefficiencies and reduced retrieval accuracy. In fact, simulation requests typically contain two critical elements: the functions to be used and the options to be set. Thus, this paper proposes using functions and options as distinct retrieval queries. However, these elements are rarely stated explicitly in simulation requests; instead, they are embedded in natural language descriptions.
To address this, an agent-driven adaptive query planning strategy is developed, as shown in Fig. 2(a), which automatically extracts function-related and option-related queries from the broader request to serve as retrieval keywords. The strategy operates in two phases: semantic recognition and keyword mapping, carried out by a retrieval agent (e.g., a general LLM). In the semantic recognition phase, the agent categorizes the request into two query types: function-related and option-related. Each function-related query is then decomposed into sub-queries, each corresponding to a potential simulation function to be used. Similarly, option-related queries are broken down into sub-queries, each linked to a potential option to be configured. This systematic separation ensures independent processing of each component within the request.
@@
To enable the retrieval agent to perform both semantic recognition and keyword mapping effectively, this paper designs a structured, general action prompt that integrates chain-of-thought prompting (CoT) [21] and few-shot prompting [22] (i.e., few-shot CoT), as depicted in Fig. 2(b). Only the few-shot examples are tool-dependent, making them modular and easily adaptable. The rest of the prompt remains general and independent of specific simulation tools.
Remark: The adaptive query planning strategy goes well beyond simple keyword extraction. It employs semantic recognition via few-shot CoT, which can translate descriptions into relevant keywords. This enables the retrieval agent to infer implicit simulation functions and options from context—even when explicit terms are missing. For example, consider the request: “For the IEEE 24-bus Reliability Test System, solve the AC optimal power flow (OPF) problem with the de-commitment of expensive generators,” which is the initial request in complex task 7 for MATPOWER (see Section V for more details). Rather than extracting generic phrases like “AC optimal power flow” or “de-commitment,” the retrieval agent—guided by few-shot CoT prompting for function mapping—directly identifies “runuopf” as the keyword, i.e., the specific MATPOWER function for OPF with de-commitment.
@@ II-B Triple-based Structure Design for Knowledge Base
This section addresses the question of which knowledge base should serve as the retrieval repository. While each power system simulation tool includes a user manual with detailed instructions, this manual is not an ideal retrieval repository. The reasons are twofold: (i) User manuals are designed for human readability rather than automated retrieval; although readable, they are unstructured and inefficient for machine-driven queries, especially when manuals primarily consist of formulas, tables, and figures. (ii) The main challenge for LLMs in generating simulation code is understanding the logical dependencies between options and functions, as many options are function-dependent. Using only the user manual for retrieval fails to capture these complex relationships effectively.
To overcome these issues, this paper proposes an additional, easy-to-construct retrieval repository: a triple-based structured option document. The following details the approach to build such a document for a given simulation tool:
  • Step 1: ChatGPT-4o is employed to parse the simulation manual and automatically extract key option entities. This process produces a list of options directly from the manual, including option names, potential default values/formats, and descriptions.
  • Step 2: The same LLM is further utilized to analyze the textual context in which the options appear, thereby determining the logical relationships between each extracted option and its associated simulation function.
  • Step 3: The outputs from Steps 1 and 2 are organized into a structured text document using a triple format, where each triple comprises: (i) the option name with its default value/format, (ii) the corresponding function dependency (linking each option and its related functions), and (iii) the description of the option along with the choices of its value.
  • Step 4: Finally, domain experts review and refine the automatically generated document to ensure that the logical associations between options and functions are accurate.
Note that the inclusion of the function dependency in the document enables retrieval for logical relationships. As will demonstrated in case studies, this supplementary repository significantly enhances retrieval efficiency and improves the accuracy of simulation code generation by preserving the logical context.
II-C Adaptability
The enhanced RAG module demonstrates high adaptability to a wide range of power system simulation tools. This flexibility is achieved by capturing the core principles of simulation coding—the use of functions and options. Specifically, the module decomposes simulation requests into distinct queries related to functions and options. This decomposition is tool-independent, thereby enabling the seamless integration of new simulation platforms. Additionally, the construction of a triple-based option document—linking options to their corresponding functions and dependencies—leverages inherent logical relationships that are common across different simulation tools. Together, these design choices underscore the module’s robust adaptability in diverse simulation environments.
@@ III Enhanced Reasoning Module
The enhanced RAG module provides LLMs with retrieval results tailored to a simulation request. However, it is still essential to strengthen the LLM’s reasoning abilities. This ensures the generation of correct simulation codes based on the retrieval results. Hence, a coding agent is required (i.e., another LLM) to write codes for simulations tasks. This agent needs to fully understand its role, assigned tasks, reasoning path, and contextual knowledge, including retrieval results, when handling simulation tasks.
To address this, this paper proposes an enhanced reasoning module, as detailed in Fig. 4. It provides structured guidance, sequential reasoning steps, and contextual knowledge to support accurate code generation by the coding agent. Details are as follows.
@@ III-B Reasoning Framework
To enable systematic, tool-independent reasoning, this paper develops a few-shot CoT framework, which breaks down the simulation task into the following universal actions:
@@ III-C Knowledge Integration
  • Dynamic Retrieval Knowledge: Complements static knowledge by incorporating real-time, request-specific details. This is achieved through the integration of the aforementioned RAG module, which retrieves relevant information on option formats, values, and function dependencies, for accurate code generation.
The integration of the RAG module into the LLM-based coding agent (from the reasoning module) is established through placeholders in the prompt. These placeholders serve as symbolic flags that are dynamically replaced by retrieval results before the prompt is sent to the coding agent. As shown in the “Retrieval Knowledge” section of Fig. 4 (b), during runtime, the enhanced RAG module retrieves simulation-specific information and substitutes these placeholders with up-to-date details on simulation functions, options, and parameters. Consequently, the coding agent generates simulation code based on both this dynamically retrieved information and the static knowledge embedded in the prompt.
Eventually, by integrating both static and dynamic knowledge, as well as the above structured reasoning framework, the coding agent is expected to generate accurate code to address the simulation request.
III-D Adaptability
The enhanced reasoning module leverages few-shot chain-of-thought prompting techniques, which embody universal principles and remain tool-agnostic. Additionally, supported by the RAG module, the reasoning module integrates two levels of knowledge: the static and dynamic knowledge. The static basic knowledge provides fundamental, coarse-grained insights—such as the basic principles of using a simulation tool. Meanwhile, dynamic retrieval knowledge supplies detailed, fine-grained information. This dual-layer approach is universally applicable across different tools. As a result, the enhanced reasoning module adapts seamlessly to a variety of simulation tasks and environments.